# Geo4D: Leveraging Video Generators for Geometric 4D Scene Reconstruction

### v1

@@ Abstract
We introduce Geo4D, a method to repurpose video diffusion models for monocular 3D reconstruction of dynamic scenes. By leveraging the strong dynamic prior captured by such video models, Geo4D can be trained using only synthetic data while generalizing well to real data in a zero-shot manner. Geo4D predicts several complementary geometric modalities, namely point, depth, and ray maps. We introduce a new multi-modal alignment algorithm to align and fuse these modalities, as well as multiple sliding windows, at inference time, thus obtaining robust and accurate 4D reconstruction of long videos. Extensive experiments across multiple benchmarks show that Geo4D significantly surpasses state-of-the-art video depth estimation methods, including recent methods such as MonST3R, which are also designed to handle dynamic scenes.
Figure 1: Geo4D repurposes a video diffusion model [102] for monocular 4D reconstruction. It uses only synthetic data for training, but generalizes well to real data. It predicts several geometric modalities, including point maps, depth maps, and ray maps, fusing and aligning them to obtain state-of-the-art dynamic reconstruction even for scene with extreme object and camera motion.
@@ 1 Introduction
4D reconstruction from videos is related to multi-view static 3D reconstruction, which is typically addressed using methods from visual geometry like bundle adjustment. New neural networks [92, 89] have recently emerged as powerful tools that can replace, or at least complement, bundle adjustment. They excel especially in difficult reconstruction scenarios, involving, e.g., textureless surfaces and occlusions, thanks to the priors that they learn from data. Because of the additional challenges involved in 4D reconstruction, we expect that such priors would benefit this task even more.
In fact, powerful networks like DUSt3R [92], designed for static multi-view 3D reconstructions, have recently been extended to the dynamic case, for example by MonST3R [113]. However, these models are heavily engineered to solve specific 3D reconstruction problems. Most importantly, they require significant amounts of training data with 3D annotations for supervision. Such data is difficult to collect for dynamic scenes, especially in real life. This suggests using 4D synthetic training data instead. However, this data is difficult to obtain at scale, and the gap with the real world can compromise generalization.
One way to mitigate this problem is to pre-train the model on tasks related to 3D reconstruction for which real data is easily available. For example, DUSt3R and derived methods use image matching for pre-training [98]. Here, we suggest starting instead from an off-the-shelf video generator. Video generators are powerful models, often considered proxies of world simulators [37, 54, 59]. More importantly for us, the videos that they generate show an understanding of effects like camera motion and perspective as well as typical object motion in the context of that scene. However, they only generate pixels, leaving any 3D or 4D understanding implicit and thus not directly actionable.
In this work, we show that a pre-trained off-the-shelf video generator can be turned into an effective monocular feed-forward 4D reconstructor. To this end, we introduce Geo4D, a novel approach for adapting Video Generators for Geometric 4D Reconstruction. With Geo4D, we demonstrate that these generic video architectures can successfully solve complex 4D reconstruction tasks, which is a step towards future video foundation models that natively integrate 4D geometry. Prior works like Marigold [28] and concurrent work DepthCrafter [22] have looked at adapting, respectively, image and video generators for depth estimation. Here, we go one step further and consider the full recovery of 4D geometry, including camera motion and dynamic 3D structure.
@@
Viewpoint-invariant point maps are a powerful representation because they implicitly encode the camera motion and intrinsics, and because they can be easily predicted by a neural network [92]. However, they are not necessarily the best representation for all parts of the scene, particularly for points far away from the observer or even at infinity, such as the sky. We thus consider two more modalities, namely disparity maps and camera ray maps, with better dynamic range. Ray maps, in particular, are defined for all image pixels regardless of the scene geometry.
Our model thus predicts three modalities: point, disparity, and ray maps. These modalities are redundant in principle, but complementary in practice. At test time, we reconcile them via a fast, global optimization step and show that this leads to significantly more robust 4D reconstructions. Due to depth and ray map prediction, we show in particular very strong empirical results on video depth estimation and in the recovery of the camera orientation.
One reason why monocular 4D reconstruction is difficult is that it is ambiguous, much more so than static 3D reconstruction. The stochastic nature of the video generator can help us deal with this ambiguity. We also introduce uncertainty maps in the encoder-decoder architecture that processes the geometric maps and that are integrated into the multi-modal alignment process.
Overall, our contributions are as follows. (i) We introduce Geo4D, a 4D feed-forward network for dynamic scene reconstruction that builds on top of an off-the-shelf video generator. (ii) We suggest generating multiple partially-redundant geometric modalities and fusing them at test time via light-weight optimization. (iii) We show the benefits of this multi-modal fusion in terms of improved 4D prediction accuracy. Our experiments show that this model is able to reconstruct even highly dynamic scenes (such as the drifting scene in DAVIS [23] presented in Fig. 1) and outperforms current video depth and camera rotation estimation.
@@ 2.1 Dynamic Scene Reconstruction
Feed-forward 3D reconstruction has achieved remarkable success across various representations, including voxels [11, 83, 74], meshes [90, 18, 72], and point clouds [41, 110]. These advancements have been further propelled by implicit neural representations [56, 60, 52, 75] and the emergence of 3D Gaussian Splatting [29, 7, 79, 9, 80, 76]. Recently, DUSt3R [92] introduced a point map representation for scene-level 3D reconstruction, followed by [86, 35, 104, 89]. However, these models predominantly focus on static 3D reconstruction. Our approach also uses point maps as a representation but extends it to handle dynamic scenes, which presents more challenges, where objects undergo motion over time.
@@
Iterative or optimization-based approaches reconstruct 4D models from monocular videos by iteratively fitting the observed data. Classical techniques often rely on RGB-D sensors [53, 24], but such steps are impractical for many real-world scenes. Recently, with advancements in neural representations [56, 52], NeRF-based approaches [57, 58, 38, 62, 39, 27] have shown impressive results. However, volume rendering in NeRF is computationally expensive. Convergence and rendering speed can be improved by using 3D Gaussian Splatting (3D-GS) representations [29, 99, 107, 91, 12, 43, 34, 111], which reduce but do not eliminate the cost of iterative optimization. Very recently, MegaSaM [40] achieved highly accurate and robust camera pose estimation and reconstruction for dynamic videos, but it requires accurate monocular depth prior. Similarly, Uni4D [108] produces accurate 4D reconstruction by leveraging various visual foundation models and performing multi-stage bundle adjustment. In contrast, our approach is a diffusion-driven feed-forward framework, which eliminates the need for per-video bundle adjustment and other depth estimation models.
@@
Similar to our approach, recent works have started to explore feed-forward 4D reconstruction for dynamic scenes: a monocular video with dynamic objects is processed by a neural network to recover a 4D representation. For objects, L4GM [66] and Animate3D [26] first generate multi-view videos from a monocular video input, and subsequently apply 3D-GS [29] to reconstruct a temporally consistent 4D model. For scenes, a notable example is MonST3R [113], which adapts the static scene reconstruction of DUSt3R [92] to handle dynamic scenes. Very recently, Easi3R [8] applies attention adaptation during inference and performs 4D reconstruction based on DUSt3R [92] in an efficient training-free manner.
2.2 Geometric Diffusion Model
Our method builds upon advancements in video diffusion models [21, 88, 73, 4, 16, 94, 19, 3, 112, 102, 31], which generate temporally consistent videos from text or image prompts. Recent studies have explored the rich 3D priors embedded within large-scale pre-trained diffusion models, employing either knowledge distillation [61, 87, 42, 96, 25, 51] or fine-tuning [45, 71, 36, 46, 20, 118, 47, 85] for 3D reconstruction and generation. While these methods have significantly advanced single-object 3D reconstruction from sparse inputs, they remain largely constrained to static, isolated objects centered within an image. Beyond single-object reconstruction, several recent efforts have extended pre-trained diffusion models to tackle scene-level 3D tasks, such as optical flow estimation [69], view synthesis [15, 68, 10, 109, 44, 81], depth estimation [13, 28, 117], and normal estimation [63, 14, 33]. More related to our approach, Matrix3D [49] jointly predicts the depth and camera parameters, and WVD [115] introduces a hybrid RGB+point maps representation for scene reconstruction. However, these approaches assume static 3D environments, whereas we address dynamic 4D scene reconstruction, which is a much harder problem due to object motion across time.
More related to our approach, concurrent GeometryCrafter [103] introduced a point map VAE with a dual encoder-decoder architecture to improve reconstruction accuracy. However, their point maps are defined in individual camera coordinates, necessitating the use of additional segmentation [30] and tracking models [101] to recover the global point map and estimate camera poses. Aether [82], on the other hand, outputs depth maps and ray maps from a video diffusion model for 4D reconstruction. In contrast, our experiments demonstrate that performance can be significantly enhanced by jointly predicting multiple geometric modalities that capture diverse dynamic ranges, ensuring better temporal coherence and robustness. Importantly, our approach is self-contained and does not rely on external models, enhancing its generality and reliability.
@@ 3 Method
Our goal is to learn a neural network $f_{\bm{\theta}}$ that can reconstruct dynamic 3D scenes from monocular videos. Given as input a monocular video =$\mathcal{I}=\{\bm{I}^{i}\}_{i=1}^{N}$ consisting of $N$ frames, where each frame is an RGB image $\bm{I}^{i}\in\mathbb{R}^{H\times W\times 3}$, the network $f_{\bm{\theta}}$ returns a representation of its 4D geometry:
@@
The network computes the disparity map $\bm{D}^{i}\in\mathbb{R}^{H\times W\times 1}$, the viewpoint-invariant point map $\bm{X}^{i}\in\mathbb{R}^{H\times W\times 3}$, and the ray map $\bm{r}^{i}\in\mathbb{R}^{H\times W\times 6}$ of each frame $\bm{I}^{i}$, $i\in 1,\dots,N$. As we discuss in Sec. 3.2, these quantities collectively represent the 4D geometry of a scene, including its dynamic structure and time-varying camera extrinsic and intrinsic parameters. No camera parameters are provided as input, so these are implicitly estimated by the model too.
We implement $f_{\bm{\theta}}$ as a video diffusion model, where ${\bm{\theta}}$ are the learnable parameters of the model. We discuss the relevant background on video diffusion models in Sec. 3.1. Then, in Sec. 3.2, we describe how we extend the model to predict the three modalities of the 4D geometry. Finally, in Sec. 3.3, we describe how we fuse and align these modalities to obtain a coherent 4D reconstruction at test time.
@@ 3.1 Preliminaries: Video Diffusion Model
Our key insight is that by building on pre-trained video diffusion models, our approach can exploit the strong motion and scene geometry priors inherently encoded within these models. Specifically, we build Geo4D on top of the DynamiCrafter [102], a “foundation” video diffusion model. DynamiCrafter is a latent diffusion model [67]: it uses a variational autoencoder (VAE) to obtain a smaller video representation and thus reduce the computational complexity. During training, a target sequence $\mathcal{X}=\bm{x}^{1:N}$, is first encoded into the latent space using the encoder $\bm{z}_{0}^{1:N}=\mathcal{E}(\bm{x}^{1:N})$, and then perturbed by $\bm{z}_{t}^{1:N}=\sqrt{\bar{\alpha_{t}}}\bm{z}_{0}^{1:N}+\sqrt{1-\bar{\alpha_{t}}}\epsilon^{1:N}$, where $\epsilon\sim\mathcal{N}(\bm{0},\bm{I})$ is the Gaussian noise, and $\bar{\alpha_{t}}$ is the noise level at step $t$ of $T$ noising steps. The denoising network $\epsilon_{\bm{\theta}}$ is then trained to reverse this noising process by optimizing the following objective:
@@ 3.2 Multi-modal Geometric 4D Diffusion
We first provide a more precise description of the 4D multi-modal representation output by our model, and then explain how the latter is encoded in the latent space for generation.
@@
The dynmic 3D structure of a scene is represented by a sequence of point maps $\{\bm{X}^{i}\}_{i=1}^{N}$, one for each of its $N$ frames. Let $(u,v)$ denote the pixel coordinates in the image plane. Then, the value $X^{i}_{uv}\in\mathbb{R}^{3}$ is the 3D coordinate of the scene point that lands at pixel $(u,v)$ in frame $\bm{I}^{i}$, expressed in the reference frame of camera $i=1$. Because the reference frame is fixed and independent of the time-varying viewpoint, we call these point maps viewpoint invariant. The advantages of this representation are convincingly demonstrated by DUSt3R [92]. For a static scene, or by knowing which image pixels correspond to the static part of a scene, knowledge of the point maps allows recovery of the intrinsic and extrinsic camera parameters as well as the scene depth. This is done by solving an optimization problem that aligns the dynamic point maps with a pinhole camera model.
As noted in Sec. 1, while point maps $\{\bm{X}^{i}\}_{i=1}^{N}$ fully encode the 4D geometry of the scene, they may not be as effective for all parts of the scene. Their dynamic range is limited, and in fact point maps are not even defined for points at infinity (sky). Hence, consider two more modalities, namely disparity maps $\{\bm{D}^{i}\}_{i=1}^{N}$ and camera ray maps $\{\bm{r}^{i}\}_{i=1}^{N}$, also encouraged by prior evidence [14, 33, 49] that diffusion model can benefit from learning to predict multiple quantities. Disparity maps are not viewpoint-invariant, but have better dynamic range than point maps (the disparity is zero for points at infinity). Ray maps only represent the camera parameters and are defined for all image pixels, independent of the scene geometry. For the disparity map, the value $D^{i}_{uv}$ is the disparity (inverse depth) of the scene point that lands at pixel $(u,v)$, as seen in frame $\bm{I}^{i}$. For the ray map, we adopt the Plücker coordinates [75, 97, 118], i.e., $\bm{r}_{uv}=(\bm{d}_{uv},\bm{m}_{uv})$, where $\bm{d}_{uv}=\mathbf{R}^{\top}\mathbf{K}^{-1}(u,v,1)^{\top}$ is the ray direction and $\bm{m}_{uv}=-\mathbf{R}^{\top}\mathbf{t}\times\bm{d}_{uv}$, where $(\mathbf{R},\mathbf{K},\mathbf{t})$ are the camera’s rotation, calibration, and translation parameters.
@@
The three modalities come in the form of images and can thus be naturally predicted by the video diffusion architecture. However, this requires first mapping them to the latent space, for which we need suitable versions of the encoder $\mathcal{E}$ and decoder $\mathcal{D}$ from Sec. 3.1. Related prior works [28, 14] for depth prediction simply repurposed a pre-trained image encoder-decoder without change. We found this to work well for depth and ray maps, but not for point maps. Hence, for the point maps only, we fine-tune the pre-trained decoder $\mathcal{D}$ using the following objective function [100]:
@@
where $\bm{\sigma}\in\mathbb{R}^{H\times W}$ is the uncertainty of the reconstructed point map, which is also predicted by an additional branch of our VAE decoder. We leave the encoder $\mathcal{E}$ unchanged to modify the latent space as little as possible; instead, we normalize the point maps in the range of $[-1,1]$ to make them more compatible with the pre-trained image encoder.
@@
The original video diffusion model is conditioned on a single image, but here we need to condition it on the entire input video =$\mathcal{I}=\{\bm{I}^{i}\}_{i=1}^{N}$. To this end, we use a hybrid conditioning mechanism using two streams.
As shown in the Fig. 2, in one stream, we extract a global representation of each frame $\bm{I}^{i}$ by passing it to CLIP [64] followed by a lightweight learnable query transformer [1]. These vectors are incorporated in the transformer via cross-attention layers injected in each U-Net block. In the other stream, we extract local spatial features from the VAE encoder and concatenate them channel-wise to the noised latents, encoding the generated 4D modalities $\{(\bm{D}^{i},\bm{X}^{i},\bm{r}^{i})\}_{i=1}^{N}$.

### v2

@@ Abstract
We introduce Geo4D, a method to repurpose video diffusion models for monocular 3D reconstruction of dynamic scenes. By leveraging the strong dynamic priors captured by large-scale pre-trained video models, Geo4D can be trained using only synthetic data while generalizing well to real data in a zero-shot manner. Geo4D predicts several complementary geometric modalities, namely point, disparity, and ray maps. We propose a new multi-modal alignment algorithm to align and fuse these modalities, as well as a sliding window approach at inference time, thus enabling robust and accurate 4D reconstruction of long videos. Extensive experiments across multiple benchmarks show that Geo4D significantly surpasses state-of-the-art video depth estimation methods.
Figure 1: Geo4D repurposes a video diffusion model [102] for monocular 4D reconstruction. It uses only synthetic data for training, yet generalizes well to out-of-domain real videos. It predicts several geometric modalities, including point maps, disparity maps, and ray maps, fusing and aligning them to obtain state-of-the-art dynamic reconstruction even for scenes with extreme object and camera motion.
@@ 1 Introduction
4D reconstruction from videos is related to multi-view static 3D reconstruction, which is typically addressed using methods from visual geometry like bundle adjustment. Recent neural networks [92, 89] have emerged as powerful tools that can replace, or at least complement, bundle adjustment. They excel especially in difficult reconstruction scenarios, involving, e.g., textureless surfaces and occlusions, thanks to the priors they learn from data. Given the additional challenges involved in 4D reconstruction, we expect that such priors would benefit this task even more.
In fact, powerful networks like DUSt3R [92], designed for static multi-view 3D reconstruction, have recently been extended to the dynamic case, for example by MonST3R [113]. However, these models are heavily engineered to solve specific 3D reconstruction problems. Most importantly, they require significant amounts of training data with 3D annotations for supervision. Such data is difficult to collect for dynamic scenes, especially in real life. This suggests using 4D synthetic training data instead. However, this data is difficult to obtain at scale, and the gap with the real world can compromise generalization.
One way to mitigate this problem is to pre-train the model on tasks related to 3D reconstruction for which real data is easily available. For example, DUSt3R [92] and derived methods [113] use image matching for pre-training [98]. Here, we suggest starting instead from an off-the-shelf video generator. Video generators are powerful models, often considered proxies of world simulators [37, 54, 59]. More importantly for us, the videos they generate demonstrate an understanding of effects like camera motion and perspective, as well as typical object motion in the context of a scene. However, they only generate pixels, leaving any 3D or 4D understanding implicit and thus not directly actionable.
In this work, we show that a pre-trained off-the-shelf video generator can be turned into an effective monocular feed-forward 4D reconstructor. To this end, we introduce Geo4D, a novel approach for adapting Video Generators for Geometric 4D Reconstruction. With Geo4D, we demonstrate that these generic video architectures can successfully solve complex 4D reconstruction tasks, which is a step towards future video foundation models that natively integrate 4D geometry. Prior work such as Marigold [28] and concurrent work DepthCrafter [22] have looked at adapting, respectively, image and video generators for depth estimation. Here, we go one step further and consider the full recovery of 4D geometry, including camera motion and dynamic 3D structure.
@@
Viewpoint-invariant point maps are a powerful representation because they implicitly encode the camera motion and intrinsics and can be easily predicted by a neural network [92]. However, they are not necessarily the best representation for all parts of the scene, particularly for points far away from the observer or even at infinity, such as the sky. We thus consider two more modalities with better dynamic range, namely disparity maps and camera ray maps. Ray maps, in particular, are defined for all image pixels regardless of the scene geometry.
Our model thus predicts three modalities: point, disparity, and ray maps. These modalities are redundant in principle, but complementary in practice. At test time, we reconcile them via a fast, global optimization step and show that this leads to significantly more robust 4D reconstructions. Due to depth and ray map prediction, we show very strong empirical results on video depth estimation and in the recovery of the camera orientation.
One of the challenges of monocular 4D reconstruction is that it is ambiguous, significantly more so than static 3D reconstruction. However, the stochastic nature of the video generator can help deal with this ambiguity. We also introduce uncertainty maps in the encoder-decoder architecture that processes the geometric maps, and integrate them into the multi-modal alignment process.
Overall, our contributions are as follows. (i) We introduce Geo4D, a 4D feed-forward network for dynamic scene reconstruction that builds on top of an off-the-shelf video generator. (ii) We suggest generating multiple partially redundant geometric modalities and fusing them at test time via lightweight optimization. (iii) We show the benefits of this multi-modal fusion in terms of improved 4D prediction accuracy. Experiments show that this model can reconstruct even highly dynamic scenes (such as the drifting scene in DAVIS [23] presented in Fig. 1) and outperforms current video depth and camera rotation estimation methods.
@@ 2.1 Dynamic Scene Reconstruction
Feed-forward 3D reconstruction has achieved remarkable success across various representations, including voxels [11, 83, 74], meshes [90, 18, 72], and point clouds [41, 110]. These advancements have been further driven by implicit neural representations [56, 60, 52, 75] and the emergence of 3D Gaussian Splatting (3D-GS) [29, 7, 79, 9, 80, 76]. Recently, DUSt3R [92] introduced a point map representation for scene-level 3D reconstruction, followed by [86, 35, 104, 89]. However, these models predominantly focus on static 3D reconstruction. Our approach also uses point maps as a representation but extends them to handle dynamic scenes, which present additional challenges due to object motion over time.
@@
Iterative or optimization-based approaches reconstruct 4D models from monocular videos by iteratively fitting the observed data. Classical techniques often rely on RGB-D sensors [53, 24], but such steps are impractical for many real-world scenes. Recently, with advancements in neural representations [56, 52], NeRF-based approaches [57, 58, 38, 62, 39, 27] have shown impressive results. However, volume rendering in NeRF is computationally expensive. Convergence and rendering speed can be improved by using 3D-GS representations [29, 99, 107, 91, 12, 43, 34, 111], which reduce but do not eliminate the cost of iterative optimization. Very recently, MegaSaM [40] achieved highly accurate and robust camera pose estimation and reconstruction for dynamic videos, but it requires accurate monocular depth priors. Similarly, Uni4D [108] produces accurate 4D reconstructions by leveraging various visual foundation models and performing multi-stage bundle adjustment. In contrast, our approach is a diffusion-driven feed-forward framework, which eliminates the need for per-video bundle adjustment and external depth estimation models.
@@
Similar to our approach, recent works have started to explore feed-forward 4D reconstruction for dynamic scenes: a monocular video with dynamic objects is processed by a neural network to recover a 4D representation. For objects, L4GM [66] and Animate3D [26] first generate multi-view videos from a monocular video input, and subsequently apply 3D-GS [29] to reconstruct a temporally consistent 4D model. For scenes, a notable example is MonST3R [113], which adapts the static scene reconstruction of DUSt3R [92] to handle dynamic scenes. Very recently, Easi3R [8] applies attention adaptation during inference and performs 4D reconstruction based on DUSt3R [92] in an efficient, training-free manner.
2.2 Geometric Diffusion Models
Our method builds upon advancements in video diffusion models [21, 88, 73, 4, 16, 94, 19, 3, 112, 102, 31], which generate temporally consistent videos from text or image prompts. Recent studies have explored the rich 3D priors embedded within large-scale pre-trained diffusion models, employing either knowledge distillation [61, 87, 42, 96, 25, 51] or fine-tuning [45, 71, 36, 46, 20, 118, 47, 85] for 3D reconstruction and generation. While these methods have significantly advanced single-object 3D reconstruction from sparse inputs, they remain largely constrained to static, isolated objects centered within an image. Beyond single-object reconstruction, several recent efforts have extended pre-trained diffusion models to tackle scene-level 3D tasks, such as optical flow estimation [69], view synthesis [15, 68, 10, 109, 44, 81], depth estimation [13, 28, 117], and normal estimation [63, 14, 33]. More related to our approach, Matrix3D [49] jointly predicts depth and camera parameters, and WVD [115] introduces a hybrid RGB+point map representation for scene reconstruction. However, these approaches assume static 3D environments, whereas we address dynamic 4D scene reconstruction, which is a much harder problem due to object motion across time.
More closely related to our approach, concurrent GeometryCrafter [103] introduced a point map VAE with a dual encoder-decoder architecture to improve reconstruction accuracy. However, their point maps are defined in individual camera coordinates, necessitating the use of additional segmentation [30] and tracking models [101] to recover the global point map and estimate camera poses. Aether [82], on the other hand, outputs depth maps and ray maps from a video diffusion model for 4D reconstruction. In contrast, our experiments demonstrate that performance can be significantly enhanced by jointly predicting multiple geometric modalities that capture diverse dynamic ranges, ensuring better temporal coherence and robustness. Importantly, our approach is self-contained and does not rely on external models, enhancing its generality and reliability.
@@ 3 Method
Our goal is to learn a neural network $f_{\bm{\theta}}$ that can reconstruct dynamic 3D scenes from monocular videos. Given as input a monocular video $\mathcal{I}=\{\bm{I}^{i}\}_{i=1}^{N}$ consisting of $N$ frames, where each frame is an RGB image $\bm{I}^{i}\in\mathbb{R}^{H\times W\times 3}$, the network $f_{\bm{\theta}}$ returns a representation of its 4D geometry:
@@
The network computes the disparity map $\bm{D}^{i}\in\mathbb{R}^{H\times W\times 1}$, the viewpoint-invariant point map $\bm{X}^{i}\in\mathbb{R}^{H\times W\times 3}$, and the ray map $\bm{r}^{i}\in\mathbb{R}^{H\times W\times 6}$ for each frame $\bm{I}^{i}$, $i=1,\dots,N$. As we discuss in Sec. 3.2, these quantities collectively represent the 4D geometry of a scene, including its dynamic structure and time-varying camera extrinsic and intrinsic parameters. No camera parameters are provided as input; these are implicitly estimated by the model as well.
We implement $f_{\bm{\theta}}$ as a video diffusion model, where $\bm{\theta}$} are the learnable parameters. We discuss the relevant background on video diffusion models in Sec. 3.1. Then, in Sec. 3.2, we describe how we extend the model to predict the three modalities of the 4D geometry. Finally, in Sec. 3.3, we describe how we fuse and align these modalities to obtain a coherent 4D reconstruction at test time.
@@ 3.1 Preliminaries: Video Diffusion Model
Our key insight is that by building on pre-trained video diffusion models, our approach can exploit the strong motion and scene geometry priors inherently encoded within these models. Specifically, we build Geo4D on top of DynamiCrafter [102], a “foundation” video diffusion model. DynamiCrafter is a latent diffusion model [67]: it uses a variational autoencoder (VAE) to obtain a more compact video representation and thus reduce computational complexity. During training, a target sequence $\mathcal{X}=\bm{x}^{1:N}$ is first encoded into the latent space using the encoder $\bm{z}_{0}^{1:N}=\mathcal{E}(\bm{x}^{1:N})$, and then perturbed by $\bm{z}_{t}^{1:N}=\sqrt{\bar{\alpha_{t}}}\bm{z}_{0}^{1:N}+\sqrt{1-\bar{\alpha_{t}}}\epsilon^{1:N}$, where $\epsilon\sim\mathcal{N}(\bm{0},\bm{I})$ is Gaussian noise, and $\bar{\alpha_{t}}$ is the noise level at step $t$ of $T$ noising steps. The denoising network $\epsilon_{\bm{\theta}}$ is then trained to reverse this noising process by optimizing the following objective:
@@ 3.2 Multi-modal Geometric 4D Diffusion
We first provide a more precise description of the 4D multi-modal representation output by our model, and then explain how it is encoded in the latent space for generation.
@@
The dynamic 3D structure of a scene is represented by a sequence of point maps $\{\bm{X}^{i}\}_{i=1}^{N}$, one for each of its $N$ frames. Let $(u,v)$ denote the pixel coordinates in the image plane. Then, the value $X^{i}_{uv}\in\mathbb{R}^{3}$ is the 3D coordinate of the scene point that lands at pixel $(u,v)$ in frame $\bm{I}^{i}$, expressed in the reference frame of camera $i=1$. Because the reference frame is fixed and independent of the time-varying viewpoint, we call these point maps viewpoint-invariant. The advantages of this representation are convincingly demonstrated by DUSt3R [92]. For a static scene, or by knowing which image pixels correspond to the static part of a scene, knowledge of the point maps allows recovery of the intrinsic and extrinsic camera parameters as well as the scene depth. This is done by solving an optimization problem that aligns the dynamic point maps with a pinhole camera model.
As noted in Sec. 1, while point maps $\{\bm{X}^{i}\}_{i=1}^{N}$ fully encode the 4D geometry of the scene, they are not effective for all parts of the scene. Their dynamic range is limited, and they are not even defined for points at infinity (e.g. sky). Hence, we consider two additional modalities: disparity maps $\{\bm{D}^{i}\}_{i=1}^{N}$ and camera ray maps $\{\bm{r}^{i}\}_{i=1}^{N}$, also encouraged by prior evidence [14, 33, 49] that diffusion models can benefit from learning to predict multiple quantities. Disparity maps are not viewpoint-invariant, but have a better dynamic range than point maps (the disparity is zero for points at infinity). Ray maps represent only the camera parameters and are defined for all image pixels, independent of the scene geometry. For the disparity map, $D^{i}_{uv}$ is the disparity (inverse depth) of the scene point that lands at pixel $(u,v)$, as seen in frame $\bm{I}^{i}$. For the ray map, we adopt Plücker coordinates [75, 97, 118], i.e., $\bm{r}_{uv}=(\bm{d}_{uv},\bm{m}_{uv})$, where $\bm{d}_{uv}=\mathbf{R}^{\top}\mathbf{K}^{-1}(u,v,1)^{\top}$ is the ray direction, and $\bm{m}_{uv}=-\mathbf{R}^{\top}\mathbf{t}\times\bm{d}_{uv}$, where $(\mathbf{R},\mathbf{K},\mathbf{t})$ are the camera’s rotation, calibration, and translation parameters.
@@
The three modalities come in the form of images and can thus be naturally predicted by the video diffusion architecture. However, this requires first mapping them to the latent space, for which we need suitable versions of the encoder $\mathcal{E}$ and decoder $\mathcal{D}$ from Sec. 3.1. Related prior work [28, 14] for depth prediction simply repurposes a pre-trained image encoder-decoder without modification. We found this to work well for disparity and ray maps, but not for point maps. Hence, for the point maps only, we fine-tune the pre-trained decoder $\mathcal{D}$ using the following objective function [100]:
@@
where $\bm{\sigma}\in\mathbb{R}^{H\times W}$ is the uncertainty of the reconstructed point map, which is also predicted by an additional branch of our VAE decoder. We leave the encoder $\mathcal{E}$ unchanged to modify the latent space as little as possible; instead, we normalize the point maps to the range [1,1][-1,1] to make them more compatible with the pre-trained image encoder.
@@
The original video diffusion model is conditioned on a single image, but here we need to condition it on the entire input video $\mathcal{I}=\{\bm{I}^{i}\}_{i=1}^{N}$. To this end, we use a hybrid conditioning mechanism with two streams.
As shown in Fig. 2, in one stream, we extract a global representation of each frame $\bm{I}^{i}$ by passing it to CLIP [64] followed by a lightweight learnable query transformer [1]. These vectors are incorporated in the transformer via cross-attention layers injected in each U-Net block. In the other stream, we extract local spatial features from the VAE encoder and concatenate them channel-wise to the noised latents, encoding the generated 4D modalities $\{(\bm{D}^{i},\bm{X}^{i},\bm{r}^{i})\}_{i=1}^{N}$.